# Deep Generative Models in Condition and Structural Health Monitoring: Opportunities, Limitations and Future Outlook

### v2

@@ 3.2.2 Aircraft structures
Aircraft SHM concentrates on ensuring structural integrity of load-bearing elements such as wings, fuselage panels, and landing gear structures. The composite wings and fuselage of commercial aircraft are subject to extreme environmental conditions from sub-zero temperatures at cruise altitudes to intense solar heating and thermal cycling on the ground. Continuous monitoring and maintenance in such environment are exceedingly difficult. Typical sensing approaches primarily rely on NDT techniques, including ultrasonic sensing and visual inspection, to detect potential damage such as delamination or fiber breakage. DGMs assist in aircraft SHM by generating realistic signals and features for scarce experimental dataset as shown in Table 6, enabling cross-modal synthesis, and producing conditioned damage progression data for downstream diagnostic and prognostic purposes [61].
@@ 3.2.3 Wind turbines
At the component level, blade inspection constitutes a key focus area within the SHM community. The turbine blades are constantly subjected to environmental stress from wind, rain, sand, and other particles, leaving them extremely susceptible to wear and tear [180]. Due to these harsh conditions, detecting blade damage using vibration or vision-based inspection becomes a critical task in wind turbine health monitoring.
@@ 3.2.4 Civil infrastructures
At the component level, subsystems of interest include bridges, tunnels, dams, and buildings. Common modalities span point sensors (strain gauges, accelerometers), distributed fiber optic sensing systems, NDT techniques (guided waves, acoustic emission), and remote imaging (UAV photography, LiDAR). DGMs complement these capabilities by enabling multimodal fusion, robust data imputation, synthetic generation of damage scenarios (such as cracks, corrosion), conditional/physics-aware simulation for digital twin calibration, domain adaptation across sites and sensor layouts, virtual sensing to estimate unmeasured fields, etc. Although UAV-based inspections are increasingly important for dams and large-span structures, vibration and strain-based data generation tasks remain prevalent across civil applications. The mapped DGM tasks and component-level applications are summarized in Table 6.
Industrial sectors Subsystems DGMs Data types Specific tasks Rotating machinery Gearboxes, Bearings VAE [181, 182, 183] Vibration; Acoustics Fault diagnosis GAN [78, 79, 80, 85, 81, 82, 83, 84, 86, 90, 87, 88, 89, 184, 185] Vibration; Acoustics; Thermal images Data generation GAN [132, 139, 143, 90, 133, 134, 87, 186] Vibration; Acoustics Domain adaptation and generalization Diffusion [95, 103, 96, 97, 98, 99, 100, 101, 187] Vibration; Acoustics Data generation Diffusion [140, 103] Vibration Domain adaptation and generalization LLM [160, 161, 162] Vibration; Text prompt Multimodal fusion LLM [160, 161, 162] Vibration; Text prompt Fault diagnosis, anomaly detection Aircrafts Blades GAN [188] Loading data Data generation Landing gears GAN [189] Landing gear data Data generation Wings, skins GAN [190] Visual images Data generation Diffusion [191, 192] Ultrasonic signals Data generation Wind turbines VAE [193, 194] SCADA GAN [115, 195] SCADA data Data generation Diffusion [196, 158] Blades VAE [197] Vibration Anomaly detection Diffusion [198] Visual images Anomaly detection Civil structures Bridges DAR [69] Strain Data generation VAE [199] Vibration Anomaly detection GAN [116, 124, 76, 117] Vibration; Strain; Visual images Data generation GAN [135] Vibration Domain adaptation and generalization Diffusion [91, 60, 200] Vibration Data generation Tunnels GAN [75] Vibration Data generation Buildings GAN [201] Visual images Data generation GAN [129, 130, 131] Vibration Domain adaptation and generalization Dams GAN [202] Visual images Anomaly detection
4 Challenges and future prospects
This section summarizes the current limitations of DGMs in CM/SHM into three main aspects, as described in Section 4.1: the need for efficient fine-tuning strategies, interpretable and reliable models, and mitigating the computational inefficiency of edge deployment. Corresponding future prospects will be discussed in Section 4.2.
4.1 Challenges and limitations
4.1.1 Fine-tuning techniques
DGMs are generally trained on large datasets during the initial pre-training stage. To adapt these pre-trained DGMs to specific tasks or domains, fine-tuning techniques are usually adopted by further training on a smaller, custom dataset. By leveraging knowledge from pre-training while learning nuances of the target domain, fine-tuning bridges the performance gap between the source and target domains. This also indicates that fine-tuning is among the most widely used strategies for applying DGMs in transfer learning and domain adaptation across diverse applications. Fine-tuning techniques for DGMs consist primarily of two branches: 1) Full fine-tuning, and 2) parameter-efficient fine-tuning (PEFT) [203]. Full fine-tuning updates all parameters of the pre-trained model while preserving its original architecture, and the size of the model after fine-tuning remains the same as the pre-trained model. However, this method requires a large storage space for DGMs and is also not achievable for deploying the fine-tuned models for downstream applications. While PEFT methods selectively update a small subset of parameters, drastically reducing computational and storage overheads [204]. PEFT methods are further categorized into:
Prefix/Prompt-Tuning [205]: Additional trainable prefix tokens added to the input or hidden layer (these prefixes are continuous pseudo tokens that do not correspond to real tokens), and only these prefix parameters are trained.
Adapter-Tuning [206]: Insert smaller neural network layers or modules into each layer of the pre-trained model. These newly inserted neural modules are called adapters. When fine-tuning downstream tasks, only these adapter parameters are trained.
Low-Rank Adaptation [207]: Learning a low-rank matrix with small parameters to approximate the parameter update of the model weight matrix, only the low-rank matrix parameters are optimized during training.
As illustrated in Fig. 9, LoRA operates by decomposing model updates into low-rank matrices, enabling efficient adaptation without altering the original parameter count.
In industrial applications, efficient fine-tuning enables pre-trained DGMs to generate domain-specific sensor data (such as vibration accelerations, infrared images, and acoustic emission signals) while maintaining their foundational understanding of complex patterns [61]. For instance, a DGM pre-trained on diverse mechanical fault datasets can be fine-tuned with LoRA or adapters based on the Hugging Face PEFT package [203] using only a small number of labeled data from a particular wind turbine gearbox. This enables the model to detect subtle anomalies on-site without requiring a complete retraining of the model from scratch, which is a promising direction for real-time condition monitoring and possible deployment in industry.
4.1.2 Explanability, interpretability, and trustworthiness
Explainable models pose another obstacle in deploying DGMs for industrial generative tasks. Since DGMs combine generative modeling with DNNs, which are often characterized by the notion of black-box models, it is important to explain the generated results and interpret internal DGM mechanisms, which are collectively termed as Explainable Generative AI (GenXAI) [208]. Note that explainability and interpretability represent distinct but related concepts, where the former tries to explain and illustrate model decisions or outputs in human-understandable terms [209], whereas the latter emphasizes understanding the internal processes governing model behavior [210]. In contrast to traditional DNN models, the scale of modern DGMs, particularly LLMs, with billions of parameters and vast training datasets introduces huge challenges for GenXAI research.
To date, different explainable skills have been developed in the realm of GenXAI, which can be primarily categorized into two types of methods: local explanation (seeks to explain the result) and global/mechanistic explanation (interprets its internal mechanism) [211]. Local approaches often adopt techniques from traditional DNN explainability, such as feature attribution methods (in this case, SHapley Additive exPlanations (SHAP) [212]). More recent local explanation techniques use LLMs themselves to yield explanations, e.g., through post-hoc natural language explanations, asking an LLM to build explanations into its generation process, or through data grounding (e.g., Retrieval-Augmented Generation (RAG) [213]). As an alternative, the second direction is for global or mechanistic explanation, which also means to interpret the entire model. Similar techniques have also been developed, such as attribution of DGM internals, interpretation of model internals, algorithmic understanding, and data influence [214]. To conclude the aforementioned content, we elaborate on the difference between local and global explanation as shown in Fig. 10 from four well-built facets: attribution, natural language, decomposing reasoning, and data grounding.
Trustworthiness remains a critical concern, particularly for LLMs. While these models are good at generating accurate predictions by leveraging statistical patterns and semantic patterns from training data, their reliability is inherently tied to the limitations of DNN models [215]. In LLMs, the models may get a good handle on grammar, syntax, and semantic relationships in language usage, but they may not necessarily map the words to real-world entities as humans do, which is crucial for ensuring the credibility and trustworthiness of the generated results. Instead, current LLMs can be prone to hallucinations, which result in the generation of content that deviates from facts or user input [216]. This discrepancy could pose a serious obstacle to practical implementation and cast doubts about the reliability of their diagnostic results.
4.1.3 Quantization and deployment
Deployment involves transferring the algorithm from a development environment into a real-world operational setting where it can start monitoring equipment. Model quantization is an important step before deploying large DGMs on edge devices (e.g., mobile devices on-site) to enable real-time monitoring. The key issue is to eliminate the heavy computational burden caused by the large number of parameters of DGMs, in particular LLMs [217]. At the very beginning, the fine-tuned DGMs need to be quantized so that 16-bit floating-point models are converted into a compact low-bit (4-bit or 8-bit) fixed-point format. This transformation needs to achieve a balance between the efficiency of fixed-point computation and the benefits of reduced memory usage. Other notable techniques include model pruning and knowledge distillation, both of which aim to reduce network size while preserving as much of the original model performance as possible [218]. The main steps involved in this process include: 1) Efficient fine-tuning of DGMs on the custom dataset to adapt to specific application scenarios; 2) Performing post-training quantization (PTQ) after fine-tuning, and the DGMs are quantized to lower accuracy, that is, a 4/8-bit fixed-point model; 3) Suitable model pruning to deploy the fine-tuned model on edge devices. Quantized LoRA (QLoRA) [219] is an efficient quantization method for LLM-based LoRA fine-tuning, as shown in Fig. 11. Following these steps, the pre-trained large DGMs can be applied to commercial-grade chips so that their generative ability can be used in real-world condition monitoring scenarios.
4.2 Future outlooks
4.2.1 Zero-shot/few-shot/multimodal data generation
DGMs, particularly LLMs, are revolutionizing zero-shot, few-shot, and multimodal data generation paradigms by enabling robust and diverse generalization in data-scarce scenarios. Zero-shot learning [146] focuses on diagnosing or predicting unseen samples, which seamlessly refers to domain generalization in Section 3.1.3. Few-shot learning, on the other hand, aims to achieve high performance with only a small number of training samples, often employing meta-learning techniques [220, 221], which is a broader approach where models learn to learn, enabling them to adapt to new tasks quickly with minimal training data. It can be foreseen that DGMs can be pre-trained on large text-sensor data pairs and then can be exploited for zero-shot or few-shot reasoning for either generative tasks or diagnostic tasks. Multimodal generation will also be greatly enhanced with DGMs, especially LLMs [31]. It is interesting to note that in Section 3.1.4, the multimodal data fusion aims for multimodal input data, which herein DGMs similarly enable effective generation of multimodal output data. In industrial systems, DGMs can generate diverse data streams, such as vibration spectra, thermal imaging, acoustic emissions, and even operational logs, which collectively indicate equipment health.
4.2.2 Physics-informed DGMs
Physics-informed neural networks (PINNs) [222] integrated with DGMs represent a promising direction for next-generation condition monitoring systems. By embedding the governing equations of the monitored system as regularization terms or loss functions, PINNs regulate the generated data to align with in-situ physical dynamics. In particular, combining physics-aware principles or rules with LLMs creates new opportunities to simulate industrial phenomena, such as variable loading conditions, thermal gradients, and mechanical stresses. To date, no published research has explored the application of physics-informed LLMs for industrial condition monitoring. Future work might focus on hybrid architectures in which PINNs guide LLMs to generate physically consistent degradation trajectories (e.g., bearing wear or crack propagation) for robust predictive maintenance. Additionally, enhancing physics-informed latent spaces by directly linking generative features to measurable physical parameters (e.g., strain, temperature) would strengthen the interpretability and reliability of generated data.
4.2.3 Reinforcement learning with DGMs
Reinforcement learning (RL) has proved highly effective in computer vision for autonomous driving and robotics. The application of RL for real-time monitoring is still in its initial stages. RL is known for its autonomous and adaptive learning features, and has been instrumental in the training and optimization of large-scale models as well as in solving goal-directed decision-making problems [223]. However, conventional RL struggles with sparse rewards, high-dimensional state spaces, and safety-critical constraints [224], but DGMs can address these challenges by modeling system dynamics, simulating environments, and generating synthetic failure samples. For instance, LLMs can learn latent representations of equipment degradation patterns from historical sensor data, allowing RL agents to train in simulated environments that capture rare fault modes or extreme operational conditions. This reduces reliance on real-world data while improving robustness to unseen anomalies. By integrating LLMs with RL agents, systems can achieve policy optimization through the way in which LLMs predict future states under different maintenance actions, enabling agents to explore strategies safely and efficiently. For partial observability issues, DGMs reconstruct missing sensor signals or infer hidden system states, enhancing RL decision-making under uncertainty. Future work may focus on hierarchical RL frameworks to coordinate long-term maintenance with short-term anomaly detection.
5 Conclusions
This work systematically reviews recent advances and emerging trends in deep generative models (DGMs) for industrial condition and structural health monitoring systems. We examine state-of-the-art DGMs, including deep autoregressive models, variational autoencoders, generative adversarial networks (GANs), diffusion-based models, and large language models (LLMs), and categorize their applications into four critical areas: (i) data generation, (ii) domain adaptation and generalization, (iii) multimodal data fusion, and (iv) fault diagnosis and anomaly detection (FD&AD) tasks. The review addresses key challenges across diverse industrial domains, such as rotating machinery, aircraft structures, wind turbines, and civil infrastructure. Challenges still exist, including the need for efficient fine-tuning strategies, enhanced model interpretability, and optimization techniques such as quantization for edge device implementation. In short, the following main conclusions can be drawn from this review.
GAN-based approaches dominate in data generation task, whereas diffusion-based models and LLMs remain in the early stages of industrial practice.
DGMs for domain generalization need to be further explored, especially under unseen scenarios.
Multimodal data fusion in LLM can be sensibly integrated with real-time monitoring data in the condition monitoring process.
For FD&AD tasks, the LLM can fuse different data streams such as vibration, acoustics, and even operational logging texts, but effective fine-tuning such as LoRA still needs to be in place.
For future perspectives, DGMs are still in their infancy for industrial condition and structural health monitoring; this review provides academic and industrial insights for both research and practical implementation.

### v3

@@ 3.2.2 Aircraft structures
Aircraft SHM focuses on maintaining the structural integrity of load-bearing elements such as wings, fuselage panels, and landing gear structures. The composite wings and fuselage structures in commercial aircraft are subject to extreme environmental conditions from sub-zero temperatures at cruise altitudes to intense solar heating and thermal cycling on the ground. Continuous monitoring and maintenance in such an environment is exceedingly difficult. Typical sensing approaches primarily rely on NDT techniques, including ultrasonic sensing and visual inspection, to detect potential damage such as delamination or fiber breakage. DGMs assist in aircraft SHM by generating realistic signals and features for scarce experimental dataset as shown in Table LABEL:table_6, enabling cross-modal data synthesis, and producing conditioned damage progression data for downstream diagnostic and prognostic purposes [YANG2025112996].
@@ 3.2.3 Wind turbines
At the component level, blade inspection constitutes a key focus area within the SHM community. The turbine blades are constantly subjected to environmental stress from wind, rain, sand, and other particles, leaving them extremely susceptible to wear and tear [KHAZAEE20221568]. Due to these harsh conditions, detecting blade damage using vibration or vision-based inspection becomes a critical task in wind turbine health monitoring.
@@ 3.2.4 Civil infrastructures
At the component level, subsystems of interest include bridges, tunnels, dams, and buildings. Common modalities span point sensors (strain gauges, accelerometers), distributed fiber optic sensing systems, NDT techniques (guided waves, acoustic emission), and remote imaging (UAV photography, LiDAR). DGMs complement these capabilities by enabling multimodal fusion, robust data imputation, synthetic generation of damage scenarios (such as cracks, corrosion), conditional/physics-aware simulation for digital twin calibration, domain adaptation across sites and sensor layouts, virtual sensing to estimate unmeasured fields, etc. Although UAV-based inspections are increasingly important for dams and large-span structures, vibration and strain-based data generation tasks remain prevalent across civil applications. The mapped DGM tasks and component-level applications are summarized in Table LABEL:table_6.
| VAE [NarayanaPichika, BOOYSE2020106612, WANG2022234] | Vibration signals; Acoustic signals | Fault diagnosis | ✓ Learns compact latent representations for effective fault diagnosis from noisy sensor data ✗ Reconstruction-based diagnosis struggles with subtle faults and generalization | ||
| GAN [WANG2025110760, ZHUANG2024111186, WANG2023106872, ZHANG2024114795, LIU2022110888, WAN2025109614, LI2022101552, LI20242958, SHEN2024109159, GUO2025110854, Sun10504387, WEN2024111663, XU2023107063, app13053136, 10018473] | Vibration signals; Acoustic signals; Thermal images | Data generation | ✓ Effectively synthesizes realistic fault signals to alleviate data imbalance ✗ Training instability and limited physical interpretability, especially with scarce or noisy data | ||
| GAN [Lou9789138, WANG2025110662, Zhuang9759507, GUO2025110854, ir202307, WANG2023455, Sun10504387, 10892083] | Vibration signals; Acoustic signals | Domain adaptation and generalization | ✓ Adversarial alignment reduces cross-domain distribution mismatch ✗ Severe domain shifts and unstable training can cause overfitting to synthetic domains | ||
| Rotating machinery | Gearboxes, Bearings | Diffusion [MUELLER2024107696, GUO2025115951, ZHAO2025109520, YU2024110343, Wang0309714, HerTerng, YANG2025110312, YI2024111481, CHEN2024110101] | Vibration signals; Acoustic signals | Data generation | ✓ Produces high-fidelity, diverse fault samples with stable training, alleviating data imbalance ✗ Slow iterative sampling and heavy computation hinder real-time use and large-scale augmentation |
| Diffusion [XU2024112396, GUO2025115951] | Vibration signals | Domain adaptation and generalization | ✓ Condition-guided diffusion aligns multi-domain distributions under unknown working conditions ✗ Domain adaptation requires careful conditioning design and high computation | ||
| LLM [TAO2025112127, PANG2024110312, qaid2024fdllmlargelanguagemodel] | Vibration signals; Text prompt | Multimodal fusion | ✓ Multimodal fusion by integrating heterogeneous sensor data, text, and expert knowledge ✗ Fusion reliability depends on prompt design and data quality | ||
| LLM [TAO2025112127, PANG2024110312, qaid2024fdllmlargelanguagemodel] | Vibration signals; Text prompt | Fault diagnosis, anomaly detection | ✓ Supports knowledge-driven fault diagnosis and anomaly detection under sparse, incomplete data ✗ Lacks physical grounding and needs careful fine-tuning for customized dataset | ||
| Blades | GAN [LEE2025112301] | Loading data | Data generation | ✓ Generates realistic blade fault signal data, mitigating scarcity of aircraft propeller damage samples ✗ Training instability and limited physics-awareness may produce unrealistic blade damage patterns | |
| Aircrafts | Landing gears | GAN [PAN2025110175] | Landing gear data | Data generation | ✓ Generates diverse landing gear load data, alleviating data scarcity ✗ GAN samples may violate physical load constraints without explicit physics-informed regularization |
| Wings, skins | GAN [10577252] | Visual images | Data generation | ✓ Augments wing and skin defect data under limited samples ✗ Synthetic data without considering material physics and damage mechanisms | |
| Diffusion [CAO2025113155, XinY101049] | Ultrasonic signals | Data generation | ✓ Generates high-fidelity composite wing and skin damage data ✗ Intensive training and slow sampling hinder scalability for large aircraft structural datasets | ||
| VAE [Liu_2023, 10153965] | ✓ Generates realistic SCADA data by capturing normal operational patterns ✗ Posterior collapse may suppress rare fault features in generated SCADA data | ||||
| SCADA | GAN [QU2020106610, CHATTERJEE2025103991] | SCADA data | Data generation | ✓ Mitigates data imbalance by generating diverse SCADA samples ✗ Training instability and mode collapse can distort temporal dependencies in generated SCADA data. | |
| Wind turbines | Diffusion [YAO2025127925, YAO2024111397] | ✓ Generates high-fidelity SCADA data, preserving temporal correlations and operational variability ✗ High computational cost and slow sampling for large-scale SCADA data augmentation | |||
| Blades | VAE [YANG2024117689] | Vibration signals | Anomaly detection | ✓ Unsupervised anomaly detection under limited labeled blade fault data ✗ Reconstruction errors may miss subtle damage and degrade under nonstationary conditions | |
| Diffusion [10697302] | Visual images | Anomaly detection | ✓ Unsupervised damage detection by capturing complex blade signal distributions ✗ Computation-intensive reconstruction and inference for real-time anomaly detection deployment | ||
| Bridges | DAR [CHEN2023116063] | Strain data | Data generation | ✓ Accurately reconstructs long-term bridge strain data by preserving thermal-induced dependencies ✗ Limited generative diversity and weak extrapolation under unseen conditions | |
| VAE [anaissi2023multi] | Vibration signals | Anomaly detection | ✓ Models normal bridge behavior for unsupervised anomaly detection under limited damage data ✗ Reconstruction-based indicators may miss localized anomalies under strong environmental variability | ||
| Civil structures | GAN [ZHENG2025119694, HOU2022111206, Xiaoming2021, GAO2023105277] | Vibration signals; Strain data; Visual images | Data generation | ✓ Generates realistic bridge sensor data for effective missing data imputation ✗ Training instability and mode collapse may distort long-term correlations and damage patterns. | |
| GAN [GE2024111236] | Vibration signals | Domain adaptation and generalization | ✓ Adversarial alignment improves cross-bridge domain adaptation under varying conditions ✗ Adaptation performance depends on physics priors and adversarial stability | ||
| Diffusion [ZHENG2025119685, SHU2025111783, wei2025generation] | Vibration signals | Data generation | ✓ Generates high-fidelity bridge vibration responses by preserving temporal correlations ✗ Computationally intensive training and slow sampling for large-scale data generation | ||
| Tunnels | GAN [LI2025112062] | Vibration signals | Data generation | ✓ Rapidly generates realistic tunnel vibration data under compressive sensing constraints ✗ Adversarial training instability may reduce fidelity for complex tunnel responses | |
| Buildings | GAN [asami2022data] | Visual images | Data generation | ✓ Generates diverse damaged-building roof images to augment scarce post-disaster datasets ✗ Synthetic damage patterns may lack generalization across diverse building types | |
| GAN [SOLEIMANIBABAKAMALI2023110404, LULECI2023110370, LULECI2023106146] | Vibration signals | Domain adaptation and generalization | ✓ Zero-shot and cross-domain translation, improving generalization across building damage states ✗ Domain shifts beyond training distributions can cause unrealistic translations | ||
| Dams | GAN [rs17203398] | Visual images | Anomaly detection | ✓ Anomaly segmentation by using infrared and visual images under complex surface conditions ✗ Detection reliability depends on training data diversity |